Four persons can hit a target correctly with probabilities respectively. If all hit at the target independently, then the probability that the target would be hit, is
Know your College Admission Chances Based on your Rank/Percentile, Category and Home State.
Get your JEE Main Personalised Report with Top Predicted Colleges in JoSA
Let persons be A, B, C, D
P (Hit) = 1 – P (none of them hits)